# Conditional Dependence via U-Statistics Pruning

### v1

@@ Abstract
The problem of measuring conditional dependence between two random phenomena arises when a third one (a confounder) has a potential influence on the amount of information shared by the original pair. A typical issue in this challenging problem is the inversion of ill-conditioned autocorrelation matrices. This paper presents a novel measure of conditional dependence based on the use of incomplete unbiased statistics of degree two, which allows to re-interpret independence as uncorrelatedness on a finite-dimensional feature space. This formulation enables to prune data according to the observations of the confounder itself, thus avoiding matrix inversions altogether. Moreover, the proposed approach is articulated as an extension of the Hilbert-Schmidt independence criterion, which becomes expressible through kernels that operate on 4-tuples of data.
@@
Hilbert-Schmidt independence criterion (HSIC), kernel methods, co-information, U-Statistics, conditional dependence.
@@ I Introduction
The problem of measuring statistical dependence emerges in many fields, such as multivariate statistical analysis [1, Ch. 6], wireless sensor networks [2, Ch. 2], data analytics [3], and statistical array processing [4]. The Hilbert-Schmidt independence criterion (HSIC[5] provides a versatile non-negative measure of statistical dependence that becomes null if and only if the data pair is independent. Grounded in the theory of reproducing kernel Hilbert space s (RKHS), the HSIC succeeds in discovering both linear and nonlinear relationships in data. A well-known interpretation of the HSIC is that dependence produces a correlation of distances [6], i.e. close couples in one data set coincide with close couples in the other data set. Another insightful interpretation is that the HSIC measures statistical dependence by gauging the correlation of data implicitly mapped onto an infinite-dimensional space. This basic idea of translating dependence to correlation also admits a more traditional implementation by explicitly mapping data onto a finite-dimensional space [7], with the advantage of leveraging classical second-order statistics for inferring information-theoretic measures.
@@
Conditional dependence becomes relevant when a third phenomenon might explain, mediate, or confound the apparent relationship exhibited by the original random pair, which happens in fields such as causal discovery and Bayesian network learning [8]. Although mapping data onto a space of higher dimension is still useful for that purpose, it is still necessary to measure conditional correlation in this newfound space. This method usually leads to the computation of the Schur complement of high-dimensional matrices [9, Ch. 4], which is known to have numerical issues concerned with matrix inversions and regularization due to the usually low rank structure of the involved autocorrelation matrices [10].
@@
This letter presents a procedure for measuring conditional dependence by statistically conditioning the observed data to a potential confounder, which bypasses the matrix inversion problems. First, a novel theoretical derivation of the classical HSIC is provided in Sec. II, by translating the problem of statistical dependence into one of correlation after mapping the data onto finite-dimensional spaces based on steering vectors [7]. Sec. III briefly reviews the theory of unbiased statistics (U-Statistics) [11], which is then employed in Sec. IV to show that conditional dependence can be accomplished by just pruning U-Statistics under the control of the confounder [12]. The obtained measure, the conditional HSIC (C-HSIC), is based on a signal processing structure resembling kernel methods, which embraces the HSIC as a particular case when the U-Statistic is complete. However, it has the distinctive feature of being articulated on 4-tuples instead of the classical processing based on pairs of data from each phenomenon.
@@ II Marginal dependence as correlation
[$\mathbf{d}(\bm{\cdot})]_{m}\triangleq\tfrac{1}{\sqrt[4]{M}}\mathrm{G}\bigl{(}\tfrac{m}{\sqrt{M}}\bigr{)}\exp\bigl{(}\mathrm{j}\pi\tfrac{m}{\sqrt{M}}\bm{\cdot}\bigr{)},$ (1)
where $m\in\{-M/2,\dots,M/2-1\}$ and $\mathrm{G}:\mathbb{R}\to\mathbb{R}$ is an even, absolutely integrable function with unit $L^{2}$-norm and maximum at $\mathrm{G}(0)$. Given two random variables $\mathsf{x}$ and $\mathsf{y}$, we define a pair of transformed random vectors [^1] Without loss of generality, and for the sake of simplicity, we let $\bm{\mathsf{u}}$ and $\bm{\mathsf{v}}$ have the same dimensionality. $\bm{\mathsf{u}}\triangleq\mathbf{d}(\mathsf{x})$ and $\bm{\mathsf{v}}\triangleq\mathbf{d}(\mathsf{y})$. Their cross-covariance matrix $\mathbf{C}_{\bm{\mathsf{u}},\bm{\mathsf{v}}}\in\mathbb{C}^{M\times M}$ is defined as $\mathbf{C}_{\bm{\mathsf{u}},\bm{\mathsf{v}}}\triangleq\operatorname{\mathrm{E}}\bigl{[}\bm{\mathsf{u}}\bm{\mathsf{v}}^{\mathrm{H}}\bigr{]}-\operatorname{\mathrm{E}}[\bm{\mathsf{u}}]\operatorname{\mathrm{E}}[\bm{\mathsf{v}}]^{\mathrm{H}}$, where $\operatorname{\mathrm{E}}[\bm{\cdot}]$ denotes the expectation operator. Then, the following implication holds [5]:
$\lim_{M\to\infty}\lVert\mathbf{C}_{\bm{\mathsf{u}},\bm{\mathsf{v}}}\rVert_{\mathrm{F}}^{2}=0\quad\iff\quad\mathsf{x}\;{\perp\!\!\!\perp}\;\mathsf{y}\;,$ (2)
@@
Consider $L$ i.i.d. samples of $\mathsf{x}$ and $\mathsf{y}$, $\{x(l),y(l)\}_{l=1,\dotsc,L}$. From them, we obtain the two transformed complex data matrices $\mathbf{U}\in\mathbb{C}^{M\times L}$ and $\mathbf{V}\in\mathbb{C}^{M\times L}$, defined as follows:
@@
The unbiased sample cross-covariance between the transformed vectors is given by
@@
where $\mathbf{P}\triangleq\mathbf{I}-\tfrac{1}{L}\mathbf{1}\mathbf{1}^{\mathrm{T}}\in\mathbb{R}^{L\times L}$. Using (2), a marginal dependence measure is given by
$\lVert\widehat{\mathbf{C}}_{\bm{\mathsf{u}},\bm{\mathsf{v}}}\rVert_{\mathrm{F}}^{2}=\operatorname{\mathrm{tr}}(\widehat{\mathbf{C}}_{\bm{\mathsf{u}},\bm{\mathsf{v}}}^{\mathrm{H}}\widehat{\mathbf{C}}_{\bm{\mathsf{u}},\bm{\mathsf{v}}})$ $=\operatorname{\mathrm{tr}}\mathopen{}\mathclose{{}\left(\tfrac{1}{(L-1)^{2}}\mathbf{V}\mathbf{P}^{\mathrm{T}}\mathbf{U}^{\mathrm{H}}\mathbf{U}\mathbf{P}\mathbf{V}^{\mathrm{H}}}\right)$
@@
where the circularity of the trace operator has been used. To see the link with the HSIC, let us examine the limit for $M\to\infty$ of the Gramm (kernel) matrices $\mathbf{K}=\lim_{M\rightarrow\infty}\mathbf{U}^{\mathrm{H}}\mathbf{U}\in\mathbb{R}^{L\times L}$ and $\mathbf{Q}=\lim_{M\rightarrow\infty}\mathbf{V}^{\mathrm{H}}\mathbf{V}\in\mathbb{R}^{L\times L}$. The elements of $\mathbf{K}$ (and analogously those of $\mathbf{Q}$) are the following:
$[\mathbf{K}]_{l,l^{\prime}}$ $=\lim_{M\to\infty}\mathbf{d}^{\mathrm{H}}(x(l))\mathbf{d}(x(l^{\prime}))=$
$\lim_{M\to\infty}$ $\sum_{m=-\frac{M}{2}}^{\frac{M}{2}-1}\tfrac{1}{\sqrt{M}}\mathrm{G}^{2}\bigl{(}\tfrac{m}{\sqrt{M}}\bigr{)}\exp\biggl{(}\mathrm{j}\frac{2\pi(x(l^{\prime})-x(l))m}{\sqrt{M}}\biggr{)}$ (6a)
$=\int_{-\infty}^{\infty}\mathrm{G}^{2}(f)\exp(\mathrm{j}2\pi(x(l^{\prime})-x(l))f)\operatorname{\mathrm{d}\!}f,$ (6b)
where (6b) is the result of looking at (6a) as a Darboux sum. Then, given (II) and taking the limit $M\to\infty$, we finally obtain the HSIC [5, Sec. 3.1]:
@@
As an alternative to RKHS theory, the result in (7) has been derived here by means of the finite-dimensional universal mapping proposed in (1), which will play a fundamental role in the extension to the conditional case proposed later on. By defining $\kappa(\bm{\cdot})$ as the inverse Fourier transform222 From the properties imposed on $\mathrm{G}^{2}(f)$, $\kappa(\bm{\cdot})$ becomes naturally an autocorrelation translation-invariant kernel, such that $\kappa(0)=1\geq|\kappa(s)|$ and $\kappa(\infty)=\kappa(-\infty)=0$. The reader is referred to [15, Sec. 1.4] where these ideas emerge in the light of Bochner’s theorem. of $\mathrm{G}^{2}(f)$, the $(l,l^{\prime})$ entries of $\mathbf{K}$ and $\mathbf{Q}$ are just the evaluation of $\kappa(\bm{\cdot})$ at a value given by the difference between two data samples:
$[\mathbf{K}]_{l,l^{\prime}}=\kappa\bigl{(}x(l^{\prime})-x(l)\bigr{)}.$ (8)
Motivated by this structure, let us introduce a more compact and general notation that will become relevant later on
$[\mathbf{S}(\mathbf{a},\mathbf{b})]_{l,l^{\prime}}\triangleq\kappa\bigl{(}[\mathbf{b}]_{l}-[\mathbf{a}]_{l^{\prime}}\bigr{)},$ (9)
being $\mathbf{S}(\mathbf{a},\mathbf{b})$ a general kernel matrix obtained from two generic vectors $\mathbf{a}$ and $\mathbf{b}$. With it, we can compactly write both kernel matrices using (9) as follows:
$\mathbf{K}=\mathbf{S}(\mathbf{x},\mathbf{x}),\quad\mathbf{Q}=\mathbf{S}\mathopen{}\mathclose{{}\left(\mathbf{y},\mathbf{y}}\right).$ (10)
In summary, the HSIC results from measuring correlation in a finite-dimensional space based on steering vectors. From that, kernel formulations arise in a second stage once dimensionality grows without bound.
@@ III Sample covariance matrix as an incomplete U-Statistic
Next we present the U-Statistic reformulation jointly with its incomplete expression. The objective is to provide an alternative covariance matrix estimation based on pairwise differences between the original samples that will serve as an introduction to the pruning of data pairs.
Consider a list containing all the unique $K_{\max}\triangleq L(L-1)/2$ pairs that can be constructed from $L$ different samples of $\mathsf{x}$ or $\mathsf{y}$. This list is ordered arbitrarily such that a single index $k$ identifies both elements of a pair:
$k\mapsto(x(\mathrm{f}_{1}(k)),x(\mathrm{f}_{2}(k))),\quad k\mapsto(y(\mathrm{f}_{1}(k)),y(\mathrm{f}_{2}(k))).$ (11)
Functions $\mathrm{f}_{1}(\bm{\cdot})$ and $\mathrm{f}_{2}(\bm{\cdot})$ return the corresponding indices of each term of the pair. Given $K\leq K_{\max}$ indices, we construct the pairwise differences
$\mathring{\mathbf{u}}_{k}\triangleq\frac{1}{\sqrt{2}}\mathopen{}\mathclose{{}\left(\mathbf{u}_{f_{1}(k)}-\mathbf{u}_{f_{2}(k)}}\right),\quad\mathring{\mathbf{v}}_{k}\triangleq\frac{1}{\sqrt{2}}\mathopen{}\mathclose{{}\left(\mathbf{v}_{f_{1}(k)}-\mathbf{v}_{f_{2}(k)}}\right),$ (12)
for $k\in\{1,\dots,K\}$, corresponding to the samples of the new zero-mean virtual sources $\mathring{\bm{\mathsf{u}}}$ and $\mathring{\bm{\mathsf{v}}}$. Their sample cross-covariance matrix $\widehat{\mathbf{C}}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}}\in\mathbb{C}^{M\times M}$ is then the following:
@@
It is worth noting that expression (13) is in fact an instance of a U-Statistic. For $K=K_{\text{max}}$, i.e. when all data pairs are taken, we obtain the equality $\widehat{\mathbf{C}}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}}=\widehat{\mathbf{C}}_{\bm{\mathsf{u}},\bm{\mathsf{v}}}$ from U-Statistics theory [11]. In contrast, for $K<K_{\text{max}}$, i.e. when (13) is an incomplete U-Statistic, although $\widehat{\mathbf{C}}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}}$ remains unbiased, its estimation variance increases due to the pruning of data. Nevertheless, $\widehat{\mathbf{C}}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}}$ is still a consistent estimate of $\mathbf{C}_{\bm{\mathsf{u}},\bm{\mathsf{v}}}$ provided that $K\rightarrow\infty$ as $L\rightarrow\infty$.
@@
A notable property of incomplete U-statistics is that their robustness against data pruning increases the larger $L$ is [12]. To briefly illustrate this property, consider taking only the $K=\mathopen{}\mathclose{{}\left\lfloor L/2}\right\rfloor$ data pairs with no indices in common. The number of remaining, unused, data pairs is equal to $L(L-1)/2-\lfloor L/2\rfloor$, which increases with $O(L^{2})$. To only use the i.i.d., nonrepeated, terms is effectively equivalent to computing $\widehat{\mathbf{C}}_{\bm{\mathsf{u}},\bm{\mathsf{v}}}$ with half of the available samples [16]. It is then safe to assume that their contribution to the overall accuracy of the sample covariance is higher than those with repeated indices. Therefore, the larger $L$ is, the higher the amount of pairs that can be pruned for some specified degradation in the estimation accuracy of the resulting sample covariance. The implication is that $K$ in the incomplete U-Statistic can be designed to grow with $O(L)$ instead of $O(L^{2})$, which provides a lot of flexibility for pruning, and will be used for choosing the number of pairs in the next section.

### v2

@@ Abstract
The problem of measuring conditional dependence between two random phenomena arises when a third one (a confounder) has a potential influence on the amount of information between them. A typical issue in this challenging problem is the inversion of ill-conditioned autocorrelation matrices. This paper presents a novel measure of conditional dependence based on the use of incomplete unbiased statistics of degree two, which allows to re-interpret independence as uncorrelatedness on a finite-dimensional feature space. This formulation enables to prune data according to observations of the confounder itself, thus avoiding matrix inversions altogether. The proposed approach is articulated as an extension of the Hilbert-Schmidt independence criterion, which becomes expressible through kernels that operate on 4-tuples of data.
@@
Hilbert-Schmidt independence criterion (HSIC), kernel methods, conditional dependence, U-Statistics.
@@ I Introduction
Conditional dependence becomes relevant when a third phenomenon might explain, mediate, or confound the apparent relationship exhibited by an original random pair. Its measurement is an important task in causal discovery and Bayesian network learning [1, 2, 3, Ch. 7], which emerge in applied fields such as earth system sciences [4] or clinical data analysis [5]. Due to its nature, empirically estimating conditional dependence is a complex problem, given that it also requires inferring conditional marginal distributions. A known approach consists in measuring the Hilbert-Schmidt norm of the normalized conditional cross-covariance operator [6], grounded in the theory of reproducing kernel Hilbert space s (RKHS). This kernel-based method is capable of measuring statistical dependence by gauging correlation of data implicitly mapped onto an infinite-dimensional space. However, it is also known to have numerical issues concerned with matrix inversions and regularization due to the usually low-rank structure of the involved correlation matrices [6]. This also leads to a fairly complex method that scales cubically with the observed data size, becoming prohibitive for large data sets.
These issues have lately been the subject of interest of numerous lines of research. A kernel-based conditional independence test is proposed in [1] by deriving the asymptotic distribution of the appropriate test statistic under the null hypothesis, which succeeds in avoiding matrix inversions, but whose complexity still increases cubically with the sample size. Other authors have proposed to approximate kernel functions by randomly selecting features [7], by providing an alternative review to the embedding process onto the RKHS [8], or by leveraging the interpretation that dependence produces correlation of distances [9], i.e. close pairs in one data set coincide with close pairs in the other [10]. Nevertheless, none of these methods addresses the aforementioned computational issues, since they all require matrix inversions. An extension of the Hilbert-Schmidt independence criterion (HSIC[11], a measure of unconditional independence, is introduced in [12] with the objective of inferring causality, a similar endeavor to this letter, although focused on joint independence rather than conditional dependence.
@@
This letter presents a procedure for measuring conditional dependence by statistically conditioning the observed data to a potential confounder in a simple way, bypassing matrix inversion problems. The basis for conditioning is to prune pairwise differences of data under the control of the confounder [13], stemming from distance-based methods [10]. Simultaneously, the estimate is constructed as an alternative derivation of the kernel-based HSIC. While these two approaches are known to be equivalent [14], we propose here an estimate that combines them, albeit for different tasks. To do so, we first provide a novel reinterpretation of the HSIC in Sec. II by translating the problem of statistical dependence into one of correlation after mapping data onto finite-dimensional spaces based on steering vectors [15]. Sec. III briefly reviews the theory of unbiased statistics (U-Statistics) [16], which is then employed in Sec. IV to show that conditional dependence can be obtained by pruning U-Statistics. The obtained measure, named conditional HSIC (C-HSIC), is based on a classical signal processing structure but linked to kernel methods, which embraces the HSIC as a particular case when the U-Statistic is complete.
@@ II Marginal dependence as correlation
[$\mathbf{d}(\bm{\cdot})]_{m}\triangleq\tfrac{1}{\sqrt[4]{M}}\mathrm{G}\bigl{(}\tfrac{m}{\sqrt{M}}\bigr{)}\mathit{e}^{\mathrm{j}\pi\frac{m}{\sqrt{M}}\bm{\cdot}},$ (1)
@@
$\lim_{M\to\infty}\lVert\mathbf{C}_{\bm{\mathsf{u}},\bm{\mathsf{v}}}\rVert_{\mathrm{F}}^{2}=0\quad\iff\quad\mathsf{x},\mathsf{y}\text{ are independent},$ (2)
where $\lVert\bm{\cdot}\rVert_{\mathrm{F}}$ denotes the Frobenius norm. Terms $\operatorname{\mathrm{E}}[\mathbf{d}(\mathsf{x})]$ and $\operatorname{\mathrm{E}}[\mathbf{d}(\mathsf{y})]$ become dense samplings of the weighted (by $\mathrm{G}(\bm{\cdot})$) marginal characteristic function of $\mathsf{x}$ and $\mathsf{y}$, respectively, in the limit of $M\to\infty$. Similarly, $\operatorname{\mathrm{E}}[\mathbf{d}(\mathsf{x})\mathbf{d}^{\mathrm{H}}(\mathsf{y})]$ becomes a dense sampling of the weighted joint characteristic function (JCF). Therefore, (2) is effectively checking the separability of the JCF at every sample point (i.e. the factorization of $\operatorname{\mathrm{E}}[\mathbf{d}(\mathsf{x})\mathbf{d}^{\mathrm{H}}(\mathsf{y})]$ as a product of expectations), which becomes equivalent to an uncorrelatedness property [17, 15], akin to the distance covariance in [10].
Consider $L$ i.i.d. samples of $\mathsf{x}$ and $\mathsf{y}$, $\{x(l),y(l)\}_{l=1,\dotsc,L}$, from which we obtain two transformed complex data matrices $\mathbf{U}\in\mathbb{C}^{M\times L}$ and $\mathbf{V}\in\mathbb{C}^{M\times L}$, defined as follows:
@@
The sample cross-covariance between the transformed vectors is given by
@@
where $\mathbf{P}\triangleq\mathbf{I}-\tfrac{1}{L}\mathbf{1}\mathbf{1}^{\mathrm{T}}\in\mathbb{R}^{L\times L}$ is the data centering matrix, a well-known projection matrix in the signal processing and kernel methods literature [18, Appx. B.7]. Using (2), a marginal dependence measure is given by
$\lVert\widehat{\mathbf{C}}_{\bm{\mathsf{u}},\bm{\mathsf{v}}}\rVert_{\mathrm{F}}^{2}=\operatorname{\mathrm{tr}}(\widehat{\mathbf{C}}_{\bm{\mathsf{u}},\bm{\mathsf{v}}}^{\mathrm{H}}\widehat{\mathbf{C}}_{\bm{\mathsf{u}},\bm{\mathsf{v}}})$ $=\operatorname{\mathrm{tr}}\mathopen{}\mathclose{{}\left(\tfrac{1}{(L-1)^{2}}\mathbf{V}\mathbf{P}^{\mathrm{H}}\mathbf{U}^{\mathrm{H}}\mathbf{U}\mathbf{P}\mathbf{V}^{\mathrm{H}}}\right)$
@@
where the circularity of the trace operator has been used. To see the link of (II) with the HSIC, let us examine the limit for $M\to\infty$ of the kernel matrices $\mathbf{K}\triangleq\lim_{M\rightarrow\infty}\mathbf{U}^{\mathrm{H}}\mathbf{U}\in\mathbb{R}^{L\times L}$ and $\mathbf{Q}\triangleq\lim_{M\rightarrow\infty}\mathbf{V}^{\mathrm{H}}\mathbf{V}\in\mathbb{R}^{L\times L}$. The elements of $\mathbf{K}$ (and analogously those of $\mathbf{Q}$) are the following:
$[\mathbf{K}]_{l,l^{\prime}}$ =${\lim_{M\to\infty}^{}}\mathbf{d}^{\mathrm{H}}(x(l))\mathbf{d}(x(l^{\prime}))$
=${\lim_{M\to\infty}^{}}\smashoperator[]{\sum_{m=-\frac{M}{2}}^{\frac{M}{2}-1}}\tfrac{1}{\sqrt{M}}\mathrm{G}^{2}\bigl{(}\tfrac{m}{\sqrt{M}}\bigr{)}\mathit{e}^{-\mathrm{j}2\pi(x(l)-x(l^{\prime}))\frac{m}{\sqrt{M}}}$ (6)
$=\int_{\mathrlap{-\infty}}^{\mathrlap{\infty}}\mathrm{G}^{2}(f)\mathit{e}^{-\mathrm{j}2\pi(x(l)-x(l^{\prime}))f}\operatorname{\mathrm{d}\!}f\triangleq\kappa\bigl{(}x(l)-x(l^{\prime})\bigr{)},$
where $\kappa(\bm{\cdot})$ is the kernel function that results from the Fourier transform333 From the properties imposed on $\mathrm{G}^{2}(f)$, $\kappa(\bm{\cdot})$ becomes naturally an autocorrelation translation-invariant kernel, such that $\kappa(0)=1\geq\lvert\kappa(s)\rvert$ and $\kappa(\infty)=\kappa(-\infty)=0$. The reader is referred to [19, Sec. 1.4] where these ideas emerge in the light of Bochner’s theorem. of $\mathrm{G}^{2}(f)$, and the integral is the limit of the Darboux sum in the second line. As a result, the entries of $\mathbf{K}$ and $\mathbf{Q}$ are just the evaluation of $\kappa(\bm{\cdot})$ at the difference between two data samples. Then, given (II) and taking the limit $M\to\infty$, we obtain the HSIC [11, Sec. 3.1]:
@@
In summary, the alternative formulation exposed above shows that the HSIC results from measuring correlation in a finite-dimensional space based on steering vectors, where the kernel formulation arises in a second stage once the dimensionality grows without bound. The interplay between kernels and cross-covariance matrices of mapped data will be used later on to condition the kernel matrices by pruning data. Beforehand, we will reformulate the estimation of cross-covariance matrices leveraging incomplete U-statistics, which will serve as the basis for data pruning.
@@ III Sample covariance matrix as an incomplete U-Statistic
Consider a list containing all the unique $K_{\max}\triangleq L(L-1)/2$ pairs that can be constructed from $L$ different samples of a random variable. This list is ordered arbitrarily such that a single index $k$ identifies both elements of a pair. Let $\mathrm{f}_{1}(k)$ and $\mathrm{f}_{2}(k)$ be two functions that return the corresponding indices of each term of the $k$th pair. Given $K\leq K_{\max}$ indices, we construct the pairwise differences
$\mathring{\mathbf{u}}_{k}\triangleq\tfrac{1}{\sqrt{2}}\mathopen{}\mathclose{{}\left(\mathbf{u}_{\mathrm{f}_{1}(k)}-\mathbf{u}_{\mathrm{f}_{2}(k)}}\right),\quad\mathring{\mathbf{v}}_{k}\triangleq\tfrac{1}{\sqrt{2}}\mathopen{}\mathclose{{}\left(\mathbf{v}_{\mathrm{f}_{1}(k)}-\mathbf{v}_{\mathrm{f}_{2}(k)}}\right),$ (8)
for $k\in\{1,\dots,K\}$, corresponding to the samples of two virtual sources $\mathring{\bm{\mathsf{u}}}$ and $\mathring{\bm{\mathsf{v}}}$. It is worth noting that these are zero-mean variables for any pair with $\mathrm{f}_{1}(k)\neq\mathrm{f}_{2}(k)$, since $\bm{\mathsf{u}}$ and $\bm{\mathsf{v}}$ are i.i.d.. Their sample cross-covariance matrix $\widehat{\mathbf{C}}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}}\in\mathbb{C}^{M\times M}$ is then the following:
@@
Equation (9) is in fact an instance of a U-Statistic. For $K=K_{\text{max}}$, i.e. when all data pairs are taken, we obtain the equality $\widehat{\mathbf{C}}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}}=\widehat{\mathbf{C}}_{\bm{\mathsf{u}},\bm{\mathsf{v}}}$ from U-Statistics theory [16]. In contrast, for $K<K_{\text{max}}$, i.e. when (9) is an incomplete U-Statistic, although $\widehat{\mathbf{C}}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}}$ remains unbiased, its estimation variance increases due to the pruning of data. Nevertheless, $\widehat{\mathbf{C}}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}}$ is still a consistent estimate of $\mathbf{C}_{\bm{\mathsf{u}},\bm{\mathsf{v}}}$ provided that $K\rightarrow\infty$ as $L\rightarrow\infty$.
@@
A notable property of incomplete U-statistics is that their robustness against data pruning increases the larger $L$ is [13]. To briefly illustrate this property, consider taking only the $K=\mathopen{}\mathclose{{}\left\lfloor L/2}\right\rfloor$ data pairs with no indices in common. The number of remaining, unused, data pairs is equal to $L(L-1)/2-\lfloor L/2\rfloor$, which increases with $O(L^{2})$. To only use the i.i.d., unique, terms is effectively equivalent to computing $\widehat{\mathbf{C}}_{\bm{\mathsf{u}},\bm{\mathsf{v}}}$ with half of the available samples [20]. It is then safe to assume that their contribution to the overall accuracy of the sample covariance is higher than those with repeated indices. Therefore, the larger $L$ is, the higher the amount of pairs that can be pruned for some specified degradation in the estimation accuracy of the resulting sample covariance. The implication is that $K$ in the incomplete U-Statistic can be designed to grow with $O(L)$ instead of $O(L^{2})$, which provides considerable flexibility for pruning, eases the overall computational complexity, and will be used for choosing the number of pairs in the next section.
Once we have determined the incomplete U-statistic formulation of the cross-covariance matrix, we are now in a position to establish the underlying rule of $\mathrm{f}_{1}(\bm{\cdot})$ and $\mathrm{f}_{2}(\bm{\cdot})$ by looking at sample pairwise differences of the confounder. Afterwards, we will be employing the pruned cross-covariance matrix to return to kernel methods in the same way as portrayed in (7).